OpenStudy (anonymous):

Determine the limit. Lim (x->infinity) (1+sin(3/x))^(x)

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} (1+\sin \frac{ 3 }{ x })^{x}\]

hartnn (hartnn):

start by substituting 1/x = y x =..? when x tends to infinity, what value does y tend to ?

OpenStudy (anonymous):

I don't quite understand how you would substitute 1/x=y? Did you mean 1/y?

hartnn (hartnn):

x = 1/y or 1/x = y same thing

OpenStudy (anonymous):

So \[ \lim_{x \rightarrow \infty} (1+\sin(3y))^{\frac{ 1 }{ y }}\] ?

hartnn (hartnn):

when x->infinity 1/x ->0 y->0 got this ?

OpenStudy (anonymous):

Oohhh, okay. But the answer is \[e ^{3}\] which I'm having trouble figuring out.

hartnn (hartnn):

ok, you know the formula lim x->0 (1+x)^(1/x) = e ?

OpenStudy (anonymous):

Nope, how do I use that to solve this?

hartnn (hartnn):

ok, in that formula you see that its required that whatever is present with the "1+", should be present in the denominator of exponent. here we have sin 3x with "1+", so we need the exponent as 1/ sin 3x

hartnn (hartnn):

so, i'll adjust the exponent as 1/y = 1/ sin 3y times sin 3y/ y got this ?

OpenStudy (anonymous):

Yea! thanks for the explanation

hartnn (hartnn):

need further explanation ?

OpenStudy (anonymous):

yes please, if that's okay

hartnn (hartnn):

sure, now we will take the limit of exponent separately got this ??